### v1

Abstract
Mixture of Experts (MoE) LLMs have recently gained attention for their ability to enhance performance by selectively engaging specialized subnetworks or “experts” for each input. However, deploying MoEs on memory-constrained devices remains challenging, particularly when generating tokens sequentially with a batch size of one, as opposed to typical high-throughput settings involving long sequences or large batches. In this work, we optimize MoE on memory-constrained devices where only a subset of expert weights fit in DRAM. We introduce a novel cache-aware routing strategy that leverages expert reuse during token generation to improve cache locality. We evaluate our approach on language modeling, MMLU, and GSM8K benchmarks and present on-device results demonstrating 2$\times$ speedups on mobile devices, offering a flexible, training-free solution to extend MoE’s applicability across real-world applications.
[Uncaptioned image][Uncaptioned image]
Figure 1: Overview of our proposed cache-aware routing method. (Left) Our proposed method enhances the cache hit-rate of existing MoEs while retaining the original model accuracy by incorporating a cache prior. (Right) Throughput of our caching method for the 4-bit and 8-bit quantized Qwen1.5-MoE models deployed on two mobile devices with 10GB and 16GB available memory and cache size of 30 and 45 experts, respectively. Our proposed Cache-Aware Routing method significantly enhances the token generation throughput compared to the baseline leveraging a Least Recently Used (LRU) Cache. https://en.wikipedia.org/wiki/Universal_Flash_Storage
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Mixture of Experts (MoE) models have emerged as a powerful approach in large language models (LLMs), offering advantages in scalability and efficiency Fedus et al. (2022b; a). By selectively activating a subset of experts for each input, MoEs handle diverse data distributions and capture complex patterns more effectively than traditional dense models. Recent models like GPT-4 (OpenAI et al., 2024), Gemini (Team et al., 2024), and Mixtral (Jiang et al., 2024) provide evidence for MoEs’ success in leveraging specialized sub-networks to achieve superior performance.
Gemini (Team et al., 2024), and Mixtral (Jiang et al., 2024) provide evidence for MoEs’ success in leveraging specialized sub-networks to achieve superior performance. Simultaneously, small language models (SLMs) with MoE architectures, such as OLMoE Muennighoff et al. (2024), Phi-3.5-MoE Abdin et al. (2024), and Qwen-MoE Qwen Team (2024), have shown promise for deployment on memory-constrained devices due to their ability to maintain high performance with fewer active parameters per token.
Despite their advantages, deploying MoE models on memory-constrained devices like smartphones and laptops presents significant challenges. One primary issue is that the parallelism techniques enhancing efficiency in server deployments of MoEs are less applicable in on-device scenarios, where tokens are generated sequentially with a batch size of one. This limitation prevents batching or parallel processing through expert sharding, leading to increased latency and inefficiencies during inference. Additionally, MoE models have a significantly larger memory footprint than dense architectures, complicating deployment on memory-constrained devices.
The sheer size of these models often surpasses the available DRAM capacity, and loading model weights directly from flash memory is detrimental to latency. To mitigate this, DRAM can be used as a cache to store a subset of expert weights, but this approach is only effective if the cache hit rate is high. Ideally, MoE models should exhibit high cache locality, where a small subset of experts is repeatedly accessed over extended sequences of tokens. Addressing these challenges requires innovative caching-friendly expert routing strategies to manage memory effectively during inference while maintaining the performance benefits of MoE architectures.
Previous works have attempted to enhance MoE model efficiency by pre-fetching expert layers using hidden states from prior layers and employing Least Recently Used (LRU) caching to store recently used experts Eliseev & Mazur (2023); Yuan et al. (2024). However, these methods rely on assumptions about expert locality and predictability, which can be unreliable. A major issue is the lack of temporal locality in the expert selection of state-of-the-art (SoTA) MoEs, leading to inefficient LRU caching due to frequent evictions and reloads, resulting in a poor cache hit rate.
In this work, we aim to utilize already-trained, high-performing MoE models by integrating cache locality priors to enhance their suitability for memory-constrained devices. Specifically, we propose methods to improve expert reuse during token generation, thereby increasing the expert cache hit rate. This approach is training-free and can be applied directly to off-the-shelf MoE models, enhancing their efficiency for on-device deployment.
To enable efficient inference on memory-constrained devices, we first analyze the temporal locality in expert selection across four SoTA MoEs, identifying factors that contribute to poor cache hit rates. We demonstrate their sensitivity to expert dropping and random expert swapping in Section 2.3. To address these issues, we propose manipulating router logits to increase the probability of selecting cached experts, eliminating the need for additional training. In summary, this work makes the following contributions:
  • We demonstrate that SoTA MoE models are not optimized for cache efficiency, exhibiting poor latency performance when an LRU cache is implemented. Additionally, we find that these models are less affected by permutations in the selection of lower-ranked experts, indicating the feasibility of lossy routing strategies.
  • We propose a method to significantly improve the cache hit rate of experts, enhancing the throughput for token generation in existing SoTA MoEs without any training. Our approach enables balancing the trade-off between cache hit rate (and thus token generation latency) and model accuracy under a fixed memory budget.
  • We report the performance of the proposed routing strategy for language modeling and present evaluation results on MMLU and GSM8K benchmarks. Surprisingly, we show that imposing some degree of routing consistency can enhance benchmark scores while making these models more on-device friendly.
  • We present on-device results of our proposed approach on mobile devices with various memory constraints, demonstrating a consistent speed-up of up to 2×2$\times$2 × compared to the regular routing with LRU caching. Due to being training-free, our approach can be deployed across a range of real-world scenarios and diverse hardware configurations
2 Background and Motivation
In this section, we provide an overview of Mixture of Experts (MoE) models, followed by a sensitivity analysis that highlights the impact of expert selection on model performance.
2.1 Preliminaries of MoE
The sparse MoE layer consists of a set of $N$ expert networks $\{E_{1},E_{2},\ldots,E_{N}\}$, and a routing network $G$ which selects a subset of these experts based on their relevance to the input tokens. Let $\mathbf{x}\in\mathbb{R}^{d}$ denote the input token. The output of each MoE layer during inference is given by:
$\mathbf{y}$ $=\sum_{i\in\mathbf{r}[{:K}]}\mathbf{w}_{i}E_{i}(\mathbf{x})$ , (1)
$\mathbf{r}$ $=\operatorname{argsort}\mathbf{w}$ , (2)
$\mathbf{w}$ $=\sigma(\mathbf{z})=\sigma(G(\mathbf{x}))$ (3)
where $\mathbf{z}=G(\mathbf{x})$ denotes the logits assigned by the router to the experts, $\sigma$ denotes the softmax operation, and $\mathbf{r}$ is an expert ranking vector based on the weights, of which only the Top-K experts are selected. Note that the weights assigned to each expert might be re-normalized in some cases after the Top-K selection in Equation 1.
In conventional MoEs, the Top-K routing mechanism to select a subset of experts is inherently dynamic and input-dependent, but can lead to significant variability in routing decisions from one token to the next. This lack of consistency poses a fundamental challenge for achieving high cache hit-rates, as it reduces the likelihood that the selected experts for the given token are already present in cache. This can be seen in Table 2, where cache lifetimes are very short, and cache miss rates are high for all models deploying a simple LRU cache policy.
2.2 MoE deployment on memory-constrained devices
Deploying MoEs on memory-constrained devices, such as mobile phones, presents unique challenges due to limited memory resources. As shown in Figure 1 (Left), these devices typically feature a combination of DRAM and flash storage, each with distinct characteristics regarding speed and capacity. DRAM provides high bandwidth but limited capacity, while flash storage offers larger capacity at the expense of slower access speeds.
Figure 2: Expert sensitivity analysis. We show the effect of dropping or replacing experts as selected by the router. The x-axis represents the expert rank (ordered by their scores) and the y-axis shows the Wikitext validation perplexity (lower values indicate better performance). The left figure illustrates the effect of dropping all experts ranked higher than $k$, whereas the right figure depicts the impact of randomly replacing the expert at rank $k$.
In the context of MoE models, not all parameters can fit into the limited DRAM available. Static parameters, such as attention weights that do not change during inference, can be stored permanently in DRAM. In contrast, given the dynamic nature of expert selection, only a subset of experts is loaded into DRAM at any given time. This subset acts as a cache, enabling faster access to frequently/previously used experts.
A high cache hit rate is important because it means experts are quickly available from DRAM, reducing the overall latency. The cache hit rate measures the proportion of selected experts that are already present in the cache at the time of selection. Let the cache set $C$ denote the indices of the experts that are currently loaded into the DRAM cache, recall that $\mathbf{r}[:K]$ are the selected experts by the routing network. The cache hit rate is then defined as:
${\operatorname{hit-rate}}=\frac{{\operatorname{count}}(\mathbf{r}[:K]\cap C)}{K}.$ (4)
The $\operatorname{miss-rate}$ is simply $1-\operatorname{hit-rate}$.
A major motivation of our work is to enhance the cache hit rate of experts without compromising the original predictive performance of the model.
To achieve this, we design various functions to update the ranking vector $\mathbf{r}$ used to select the Top-K experts without making any other changes to the MoE forward pass as shown in Equation 1. This allows for a model-agnostic method that can be universally applied.
Our goal is to find new ranking vectors promoting experts that are currently in cache, while minimizing the effect on downstream model performance.
2.3 MoEs are less sensitive to dropping experts with lower scores
The potential for MoE models to become more cache-friendly without sacrificing task performance hinges on their robustness to variations in expert selection. To evaluate this assumption, we conduct a series of analyses on four different MoE architectures: DeepSeek-V2-Lite DeepSeek-AI et al. (2024), Qwen1.5-MoE-A2.7B Qwen Team (2024), Phi-3.5-MoE Abdin et al. (2024), and Mixtral-8x7B Jiang et al. (2024). Our findings demonstrate that these models can tolerate some deviations from the router-selected experts as long as the experts with the highest weights are kept.
The left plot in Figure 2 illustrates the impact of completely removing (pruning) experts ranked at or above a specified index. Notably, the removal of the second highest ranked expert leads to a significant performance drop across all four models. However, for models with a larger number of active experts, such as Qwen and DeepSeek, performance quickly recovers, allowing for the pruning of higher-ranked experts with minimal performance loss.
To illustrate the significance of expert rank while maintaining the total number of activated experts, we replaced the expert ranked $k$ with a randomly selected expert, as shown in the right plot of Figure 2. The results indicate that substituting the most important expert (rank one), which has the highest score, drastically affects model performance. Replacing the second ranked expert also leads to severe degradation, but for the higher ranks, the effect of swapping experts is minimal.
This sensitivity analysis of MoEs shows that dropping experts from the Top-K selection can adversely affect model accuracy and efficiency. In particular, all models are sensitive to dropping of the Top-1 and Top-2 experts. However, the results also suggest that while retaining the highest-ranked experts is crucial, most MoE models are relatively robust to minor variations in the experts chosen by the router.
3 Cache-aware expert routing
In this section, we present our primary techniques designed to enhance the consistency of routing decisions during token generation, which in turn improves the expert cache hit rate and overall model throughput. We propose a general training-free method that can be applied to off-the-shelf MoE models to improve their efficiency for on-device deployment. Additionally, we introduce two simple baselines to increase cache hit rates: Max Rank and Cumulative Probability Threshold.
Figure 3: Our proposed cache-aware routing method. A bias is added to the logits to promote the selection of in-cache experts.
3.1 Max Rank Approach
In Section 2.3 we demonstrated that MoEs exhibit some robustness to the swapping of their experts with random ones. Building on this insight, we aim to deviate from the Top-K set of experts $\mathbf{r}[:K]$, selected by the router and instead promote experts that are already present in the cache $C$. We first define a general promotion operation that elevates the ranking of all experts in an ordered subset $\mathbf{r}^{\text{subset}}$:
$\operatorname{\operatorname{promote}}(\mathbf{r}^{\text{subset}};\mathbf{r}^{\text{all}}):=\mathbf{r}^{\text{subset}}\oplus(\mathbf{r}^{\text{all}}\setminus\mathbf{r}^{\text{subset}}).$ (5)
Here $\oplus$ denotes the concatenation operation and the set subtraction $\setminus$ preserves the order of it’s left operand.
A naive solution would be to simply promote all experts currently in the cache, but this approach does not take into account the expert probability assigned by the router. Instead, we limit the promotion of elements in the cache by a maximum allowed rank $M$ such that the experts with the lowest router probabilities are not chosen:
$\mathbf{r}^{\text{max-rank}}:=\operatorname{\operatorname{promote}}(\mathbf{r}[:M]\cap C;\mathbf{r}).$ (6)
As observed in Section 2.3, certain Top-J experts are crucial for model performance. To ensure these experts are not overwritten by others in the cache, we always select them, even if they are not in the cache. This can be achieved by adding a second promotion operation:
$\operatorname{\operatorname{promote}}\Big{(}\mathbf{r}[:J];\operatorname{\operatorname{promote}}(\mathbf{r}[:M]\cap C;\mathbf{r})\Big{)}.$ (7)
Finally, the pseudocode for the max-rank algorithm with always keeping the Top-J experts is shown in Algorithm 1.
Algorithm 1 Max Rank
1:max-rank $M$, minimal-rank $J$, and original ranking $\mathbf{r}$
[…]
4:return $\mathbf{r}^{\prime}$
In practice, we apply the Top-J promotion to all of our approaches.
3.2 Cumulative Probability Threshold Approach
A limitation of the max-rank routing strategy is that it does not take into account the distribution of all expert probabilities $G(\mathbf{x})$. For instance, an input where the most probable expert has a very high probability will likely require a very low max-rank $M$ to maintain good model performance. Conversely, for an input with uniformly distributed router probabilities, there is no strong expert preference, allowing for a very high max-rank $M$ for better cache hit rates.
We address this issue in the cumulative probability threshold approach by dynamically choosing the max-rank $M$ for every layer and every input $\mathbf{x}$. We determine $M$ by summing the sorted probabilities of the router outputs $G(\mathbf{x})$ from highest to lowest until a predefined cumulative probability threshold $p$ is reached:
$M=\operatorname{min}i\ \ s.t.\sum_{j=1}^{i}G(\mathbf{x})[r_{i}]\geq p.$ (8)
Using the set of $M$ values found for each layer and each input, we apply the max-rank strategy from Equation 7 to select a new set of experts. The pseudocode for the cumulative probability threshold approach can be found in Algorithm 2. This method is more dynamic as it accounts for the uncertainty in the router’s predictions. When the set of experts required to reach the threshold is larger, it indicates less confidence from the router, allowing for more flexibility in replacing a non-cached expert with the highest probability expert available within this set that is already in the cache.
Algorithm 2 Cumsum Threshold
1:probability threshold $p$, minimal-rank $J$, original ranking $\mathbf{r}$, and expert weights $\mathbf{w}$
2:$p_{\text{cum}}\leftarrow 0$
3:$M\leftarrow 0$
4:while $p_{\text{cum}}<p$ do
5:     $M\leftarrow M+1$
6:     $p_{\text{cum}}\leftarrow p_{\text{cum}}+\mathbf{w}[r[M]]$
7:end while
[…]
10:return $\mathbf{r}^{\prime}$
3.3 Cache Prior Reranking
Although the cumulative threshold routing strategy dynamically adjusts the max-rank, it still imposes a hard limit beyond which experts are not promoted. This may be suboptimal, especially for router distributions with long tails, which may be better suited for making cache-friendly decisions.
To address this, we use a cache prior that directly manipulates the router logits $\mathbf{z}=G(\mathbf{x})$ to increase the probability of selecting experts already present in cache. Importantly, we use the manipulated logits $\mathbf{z}^{\prime}$ only to find a new ranking vector $\mathbf{r}^{\prime}$ as defined in Equation 2. We still use the unmodified router logits to compute the expert weights used in Equation 1. Figure 3 provides an overview of our proposed cache-aware routing method. Let the bitmask $\mathbf{m}_{t}\in\{0,1\}^{N}$ denote the state of the cache following the generation of the $(t-1)$-th token.
As with our other methods, we aim to ensure that the Top-J experts are always selected, regardless of wether they are in the cache. We achieve this by adding them to the bitmask $\mathbf{m}_{t}$, resulting in an updated bitmask denoted as $\tilde{\mathbf{m}}_{t}$. We then boost the logits corresponding to the bitmask as follows:
$\mathbf{z^{\prime}}=\mathbf{z}+\lambda\cdot\Delta_{\text{avg}}\cdot\tilde{\mathbf{m}}_{t},$ (9)
where $\lambda\in[0,1]$ is a scaling factor that determines the influence of the cache state on the logits. Here, $\Delta(\mathbf{z})$ is defined as:
$\Delta_{\text{avg}}=\mathop{\mathbb{E}}_{\mathbf{x}\in\mathcal{X}}\mathop{\mathbb{E}}_{t\in 1...T}[\max(\mathbf{z})-\min(\mathbf{z}))].$ (10)
which represents the range in logits for this layer that we estimate using a running average over datapoints and over tokens.
In essence, this method involves adding a fraction of $\Delta(\mathbf{z})$ to the logits of experts that are already in the cache. By utilizing the true range of weights, our cache prior becomes adaptive, allowing a single hyperparameter to yield distinct effects across different layers and tokens.
We provide an ablation on the choice of $\Delta$ in Section 4.6. Additionally, we explore the possibility of directly learning $\Delta$ in Appendix, Section D.
4 Experiments
Model Params Experts
Total Active Shared Total Top-k Params
Mixtral-8x7B 46.7B 13B 8 2 176M
Phi-3.5-MoE 41.9B 6.6B 16 2 79M
DeepSeek-V2 15.92B 2.8B 2 (64+2) (6+2) 8.6M
Qwen1.5-MoE 14.3B 2.7B 4 (60+4) (4+4) 8.6M
Table 1: The MoE architectures used in our experiments. Column Params in Experts indicates the amount of parameters per expert.
Figure 4: Trade-off curves between Wikitext perplexity and cache miss rate for four MoE models with a cache size set to half the total number of experts.
Figure 5: MMLU (5 shots) task evaluation.
Figure 6: GSM8K (8 shots) task evaluation.
Figure 7: Expert selection during token generation on a GSM8K sample (green: cache hit, red: cache miss, blue: in cache).
In this section, we evaluate the effectiveness of our cache-aware routing method (dubbed Prior for simplicity) and compare it to our Pruning, Max-Rank, and Cumulative Sum Thresholding baselines described in Section 3. We implement all methods on four SoTA MoE models: DeepSeek-V2-Lite DeepSeek-AI et al. (2024), Qwen1.5-MoE-A2.7B Qwen Team (2024), Phi-3.5-MoE Abdin et al. (2024), and Mixtral-8x7B Jiang et al. (2024). The architectural details of these models are summarized in Table 1.
The MoE layers in DeepSeek-V2-Lite include 2 shared experts activated for every token and 64 selectable experts, out of which 6 are activated per token, resulting in 2.4B active parameters per token out of a total of 15.7B parameters. Similarly, Qwen1.5-MoE-A2.7B has 4 shared experts and 60 selectable experts, with 4 activated for each token, leading to 2.7B active parameters (total 14.3B parameters). Phi-3.5-MoE comprises 16 experts using Top-2 routing, resulting in 6.6B active parameters (total 42B parameters). Lastly, Mixtral-8x7B includes 8 experts with Top-2 routing, resulting in 12.9B active parameters (total 46.7B parameters).
To assess the performance of our method, we conduct experiments three tasks: language modeling, Massive Multitask Language Understanding (MMLU), and Grade School Math (GSM8K). For language modeling, we report perplexity and cache miss rate on the validation subset of the WikiText-2-raw-v1 dataset Merity et al. (2017). Next, we evaluate downstream performance on the MMLU dataset Hendrycks et al. (2021) and the GSM8K dataset Cobbe et al. (2021). The MMLU dataset consists of multiple-choice questions spanning 57 subjects, while GSM8K evaluates multi-step reasoning on math problems. For both datasets, we report accuracy and cache miss rate.
In our experiments, we examine trade-offs between model accuracy and cache miss rate under varying DRAM cache size limits. For each method, we sweep over hyperparameters to create a pareto front of performance vs. cache miss rate.
In addition to the baseline results, each plot includes the performance of the model utilizing the Least Recently Used (LRU) caching policy. All of our routing methods implement an LRU cache, but the LRU baseline specifically refers to the model that employs the LRU cache without any re-ranking strategy. As a result, this baseline maintains accuracy without any trade-offs. In contrast, our other methods incorporate routing strategies that balance accuracy with LRU friendliness, allowing for improved cache utilization while still aiming to optimize performance.
4.1 Implementation Details
Keeping Top-J Experts
As observed in Section 2.3, certain Top-J experts are crucial for model performance. To prevent these experts from being overwritten by others in the cache, we always select them, even if they are not in the cache. In the cases of Phi and Mixtral, where only two experts are activated, we use $J=1$. For Qwen and DeepSeek, where the first major perplexity drop occurs at $J=2$, we select $J=2$.
We demonstrate the impact of the specific choice of Top-J in Figure 4. As shown, retaining the Top-J experts is crucial for the performance of baseline methods. Simultaneously, impact to our cache-aware routing method is less clear. However, its effect on our cache-aware routing method is less pronounced. To ensure a fair comparison and simplify the approach, we use the same Top-J configuration across all methods. Further discussion is presented in Section 4.6.
Dataset-Specific Details
[…]
Pruning
In our pruning baseline, we select a rank $h\leq k$ and discard all experts with a rank greater than or equal to $h$. Note that we still use $k$ to normalize the cache miss rate, as indicated in Equation 4.
LRU
For all methods in this work, when the cache is full, we rely on the LRU eviction policy to drop experts and free up space. Unlike traditional paging, there is no inherent order for incoming elements in each MoE layer, as an expert’s importance for the current token does not necessarily indicate its future relevance. We choose to add elements to the LRU stack based on their router weights, prioritizing experts with the highest weights for addition to (and subsequent eviction from) the cache. We observed that this approach yielded slightly better performance compared to the inverse method, as shown in Figure 10 in Appendix C.
4.2 Language Modeling Evaluation
In this section, we assess the effect of our proposed cache-aware routing method on the language modeling capabilities of MoE models. Figure 4 presents trade-off curves for perplexity versus cache miss rate of our Prior approach compared to the Pruning, Max-Rank, and Cumsum-threshold methods with a cache size equal to half the number of experts. The results show that our approach consistently outperforms all baselines across all ranges of perplexity and cache miss rate. Among the baseline methods, Pruning performs the worst, indicating that cache-informed expert replacement is essential for maintaining accuracy. Max-Rank consistently surpasses Pruning, while Cumsum-Threshold consistently outperforms Max-Rank. Ultimately, our Prior method Pareto-dominates all other approaches.
We also evaluate language modeling performance with only a quarter of experts cached, as shown in Figure 12 in Appendix, Section B. Our method’s robustness to different cache sizes makes it appealing for deployment across a range of devices.
4.3 Downstream Tasks Evaluation
Figure 5 and Figure 6 present results on the MMLU and GSM8K benchmarks, respectively. Our Prior method’s Pareto curve consistently outperforms other methods, often showing significant cache miss rate reductions with no accuracy loss compared to baseline LRU. Increasing the cache bias $\lambda$ improves cache miss rates until a minimum is reached due to the Top-J selection guarantee. We also observe that GSM8K has noisier accuracy results, as can be expected for a generative task. Importantly, while accuracy may fluctuate, adjustments to the hyperparameters for each routing method —namely $M$, $p$, and $\lambda$— consistently leads to a predictable effect on cache miss rates.
Interestingly, we observe that improved cache consistency (reduced miss rate) can lead to increased accuracy on both benchmarks, potentially due to a regularization effect from consistent expert routing.
Figure 7 shows a qualitative visualization of cache hits/misses for a GSM8K sample, comparing LRU and Prior($\lambda$=0.5) methods, with cache miss rates of $\sim$23and $\sim$15%, respectively, for Phi-3.5-MoE-instruct. It is evident that the Prior method results in fewer cache misses, and the cache lifetimes for this method are longer than those of the LRU baseline. These observations are further supported by Table 2, which presents the average cache miss rate and cache lifetime for the Wikitext dataset.
Model Cache Size Strategy Lifetime Miss Rate (%)
Qwen1.5-MoE 30 / 60 LRU 22 35
Prior 111 7
DeepSeek-V2 32 / 64 LRU 20 28
Prior 188 3
Phi-3.5-MoE 8 / 16 LRU 22 22
Prior 55 9
Mixtral-8x7B 4 / 8 LRU 5 40
Prior 10 21
Table 2: Cache sizes, average cache lifetimes, and miss rates for LRU and prior caching strategies with $\lambda=0.5$ on the Wikitext validation set.
4.4 On-Device Deployment
To evaluate the effectiveness of our cache-aware routing technique in real-world scenarios, we deployed the Qwen1.5-MoE-A2.7B model with our cache-aware routing on two mobile devices (with 12GB and 16GB RAM) equipped with some of the latest Qualcomm Snapdragon®  [^1] Snapdragon branded products are products of Qualcomm Technologies, Inc. and/or its subsidiaries. processors running Android 14. Using llama-cpp Gerganov (2023) for CPU-based deployment, we modified the implementation to add both the LRU caching for experts and our Prior algorithm. Additionally, we enabled memory locking (mlock) to prevent the Android OS from offloading expert weights from memory, which would otherwise require reloading for each token.
We tested our approach in two settings: (1) deploying a 4-bit quantized model on a device with 12GB RAM and (2) deploying an 8-bit quantized model on a device with 16GB RAM. In both cases, our method was applied only during autoregressive generation, allowing the model to benefit from optimizations designed for prompt processing.
In the first setting, we reserved 2GB of memory on the 12GB device, limiting available memory to 10GB (shared with the OS). Here, the cache size was limited to 45 experts per layer (out of 60). In the second setting, the 16GB device hosted the 8-bit quantized Qwen Model with a cache size limit of 30 experts per layer.
To determine optimal cache sizes for each setting, we initially deployed the model on both devices using the LRU caching policy across various cache sizes. Results, shown in Figure 9, indicate that exceeding a certain cache size reduces available memory, causing the OS to reload uncached components (e.g., kv-cache, activations) for each token, which significantly slows down inference. We used the highest LRU throughput as a reference point (1x) to demonstrate the speedup from our method independently of wall clock variations due to implementation specifics.
The quantitative results for both settings are presented in Figure 1 (Right). Box plots, computed over 10 runs per experiment, show median, minimum, and maximum values. Our approach provides over a 2×2$\times$2 × speedup on-device. Additionally, qualitative results in Table 3 in the Appendix demonstrate that this speedup has minimal impact on the quality of generated text.
4.5 Temporal Consistency of Expert Selections
In Table 2, we assess the impact of our method on average cache lifetime, defined as the average number of time steps (tokens generated) an expert remains in memory before being offloaded. Higher cache lifetime indicates better memory efficiency. We calculated this measure over the entire WikiText validation set, averaged across tokens.
Our approach significantly extends cache lifetimes. For example, using LRU, each expert in both Qwen and DeepSeek models stays in memory for only  22 tokens, often less than a few seconds on modern devices. With our method, experts remain in memory for 59×5-9$\times$5 - 9 × longer, reducing the frequency of weight loading from Flash and improving throughput.
4.6 Ablation Experiments
In this section, we perform ablations using the DeepSeek model with a cache size of 32, evaluated on Wikitext for the perplexity/cache miss rate trade-off.
Varying the Number of Top-J Experts
As discussed in Section 4.1, we always select Top-J experts ($J<K$) even if they are not in cache. Figure 4 shows the effect of varying $J$ on model performance. While baseline Max-rank and Cumsum-threshold methods clearly benefit from keeping Top-J intact, impact on our Prior method is less pronounced. It leads to improved perplexity for certain models (QWEN, Phi) while being detrimental to other models (DeepSeek, Mixtral). Since keeping $J>0$ is critical to baseline methods while providing good-enough trade-off between perplexity and cache miss rate for our Prior method, we decide to keep it fixed per-model across all methods.
It is important, however, to note that different values of $J$ optimize different constraints. For example, $J=0$ enables very low ($<2\%$) cache miss rates while, potentially, damaging language model performance. In practice, flexibility of our training-free approach allows $J$ to be adapted easily for deployment on different targets. Therefore, one can choose $J$ depending on the specifics of the target setting.
Logit Range Estimation
In our method, the scaling factor $\lambda$ is multiplied by a running average of the logits range $\Delta_{\text{avg}}$, as shown in Equation 9. This estimate of the logit range is dynamic, depending on model and input. To improve the estimate, we could alternatively calculate $\Delta_{\text{avg}}$ over a calibration dataset. We tested this by estimating the range over the entire Wikitext training subset. Results in Figure 8 show that our running average performs comparably to estimates based on the full dataset, while also offering robustness to out-of-domain data where logit ranges may differ from general text data (e.g., for a coding task).
Figure 8: Performance on WikiText for different logits estimation strategies
5 Related Work
Speculative Routing and Weight Preloading: Recently, various approaches have been developed to optimize MoE inference through speculative routing and weight preloading Cui et al. (2023); Xue et al. (2024); Eliseev & Mazur (2023); Du et al. (2024). These methods primarily focus on expert prefetching strategies to boost cache hit rates. For example, MoE-Infinity Xue et al. (2024) uses look-ahead routing and caches the most frequently used experts. SiDA-MoE employs an offline-trained hash function to reduce expert selection overhead. However, SiDA-MoE’s approximate caching results in a significant drop in perplexity, underscoring the difficulty of maintaining accuracy. Additionally, Eliseev & Mazur (2023) combines lookahead routing and LRU caching, necessitating the pre-loading of multiple experts well in advance to optimize the overlap between data movement and inference calculations. While this approach helps conceal memory transfer costs, it can adversely affect model accuracy. Speculative routing to improve cache hit-rate without loss of accuracy is particularly challenging on memory-constrained devices where token generation is more memory-bound than compute-bound. In contrast, our method enhances routing consistency and exploits locality priors, achieving minimal changes in perplexity and significant latency improvements. Our approach is orthogonal to speculative routing methods.
Efficient LLM Inference: Techniques such as LLM quantization Dettmers et al. (2022); Frantar et al. (2023); Shao et al. (2024) and compression Frantar & Alistarh (2023) can significantly reduce the memory footprint of large language models (LLMs). Recent studies have explored pruning or merging MoE experts based on their utilization to conserve memory Chen et al. (2022); Li et al. (2024); Lu et al. (2024), though this can result in decreased performance. To facilitate the deployment of off-the-shelf MoEs on devices with limited memory without a model surgery that would change the architecture, previous works have focused on optimizations like offloading parameters to host memory Alizadeh et al. (2024); Yu et al. (2024); Jung et al. (2023); Rasley et al. (2020). These approaches employ dynamic scheduling strategies to offload sparse parameters while maximizing the overlap between expert loading and computation. For instance, DeepUM Jung et al. (2023) records CUDA kernel execution patterns to predict which kernel is likely to be executed next. However, MoE model execution patterns can vary significantly depending on the task and context. Our proposed method enhances expert cache hit rates without requiring complex hardware implementations and remains agnostic to the specific task being addressed.
6 Conclusion
In this work, we introduced a novel training-free cache-aware expert routing approach for Mixture of Experts (MoE) models, tailored to enhance inference efficiency on memory-constrained hardware. By increasing cache hit rates through consistent routing, our method achieves a significant reduction in latency while preserving model accuracy, demonstrating up to a 2$\times$ speedup over regular routing with LRU caching techniques on mobile devices.
Our method is highly adaptable, requiring no model retraining, making it suitable for a wide range of hardware configurations. This adaptability reduces deployment costs and facilitates the practical application of MoE models in real-world, resource-limited environments.
Appendix
Appendix A LRU Throughput
In Figure 9 we show on-device performance for various cache sizes. Similar to Section 4.4, we use best LRU performance as a reference point of 1×1$\times$1 ×.
One can clearly see that increasing cache size beyond $30$ in case of 4-bit quantized model on 12GB device (left) and beyond $45$ in case of 8-bit quantized model on 16GB device (right) leads to decrease in performance. This is due to the fact that increasing cache size reduces available memory, causing the OS to reload uncached components (e.g., kv-cache, activations) for each token, which significantly slows down inference.
Given these results, we use same cache size of $30$ for 4B and $45$ for 8B in experiments in Section 4.4.
Figure 9: Throughput of the 4-bit and 8-bit quantized Qwen1.5-MoE models with LRU caching deployed on two mobile devices with 12GB and 16GB available memory, respectively.
Appendix B Smaller cache sizes
Throughout all our experiments in the paper, we use a cache size that is $\frac{1}{2}$ times the total number of experts. In Figure 12 we demonstrate performance of our approach for models with cache size equal to $\frac{1}{4}$ of the total amount. Our Prior method is outperforming all other methods across the board. Note that it is doing so without any changes to method itself making it appealing for final user who might want to deploy it across a range of devices.
Appendix C LRU Ablation
Figure 10: Wikitext Perplexity results for varying LRU addition approaches.
Figure 11: Prior vs Learned Prior performance.
In Figure 10 we show that our approach (Top-Bottom) for applying the LRU cache, as described in Section 4.6 performs better than Bottom-Top approach. However, the difference is fairly insignificant and thus the LRU strategy should not be considered an important factor.
Appendix D Learned Prior
In Section 3.3 we present our training-free Cache-Aware routing method. As shown throughout the paper, the presented method is very strong and competitive across a range of scenarios.
However, it is also possible to make the proposed method learnable. To do so, we propose learning the additive prior term using a small cache MLP layer conditioned on the cache state. Specifically, we implement this by employing a two-layer MLP that takes both the cache state and router logits as input, outputting a single bias vector. This bias is then added to the router logits and used for the next step. The loss function is optimized on the softmax outputs. We force weights assigned to experts that are in the cache but not in Top-K to move closer to one while penalizing weights for those that are in Top-K but not in the cache to move closer to zero.
A downside of learnable priors is that we are required to train many independent cache MLPs to obtain trade-off points between cache hit rate and model accuracy. We conducted an experiment to verify the effectiveness of learning this bias but did not observe notable improvements. This optimization is challenging and the results were close to the prior performance without outperforming it. Moreover, this approach would be less beneficial in settings with multiple deployment targets. Therefore, the training-free nature of our approach proves very advantageous for real-world applications.
Given the increased training expense, comparable performance and limited applicability, we skip this approach for now, leaving it for future work.
Figure 12: Wikitext Perplexity results for cache size that is one quarter the amount of experts.
Model Q4, Cache Size = 45 Q8, Cache Size = 30
Prompt Write a short poem inspired by “Numb” by Linkin Park but about Mixture-of-Expert models. Building a website can be done in 10 simple steps:
Step 1:
LRU The model is a type of artificial intelligence that uses multiple machine learning algorithms to make predictions or decisions. It’s designed to make better decisions faster than an expert, but it can also make mistakes due to the complexity of combining the decisions of different algorithms. In the realm of data and code, Lies a model so bold and brave, A mixture of experts, it’s its name, Working together, making decisions, it thrives. With algorithms galore, it’s a sight to behold, Predicting and deciding, with speed and force, Majestic in its complexity, A marvel of modern science. Get a website domain name. This is the address where your website will live. You can choose a name that describes your business or brand. It’s best to choose a name that’s easy for people to remember and spell.
Step 2: Find a web hosting service. A web hosting service is where your website files will be stored. There are many different types of web hosting services to choose from, including shared hosting, VPS hosting, and dedicated hosting. You’ll want to choose a hosting
[…]
Prior, $\lambda=0.8$ The first part of the English language Wikipedia choose Mixture-of-Expert models as the 2018 word. Mixture-of-Expert models use multiple neural networks as expert solutions. I’m not a poet, but here’s my attempt at a very loose and free poetry style translation of the given meaning.
A model so complex, Experts gather from near and far The experts of yore, to bring their expertise, their might A model to optimize, to help us all survive From AI to the rescue, it’s transformation we’ll rejoice
In the mind of Chester and Patel, Expert insights to bestow, Insights to empower,
Find a Domain Name Before you start building your website, you will need to choose a domain name (the name of your website). Your domain name is the address that people will use to reach your website, so you want to make sure it’s easy to remember and spell and related to what your website is all about. For example, if you’re building a website for your dog grooming business, your domain name could be.
Step 2: Choose a Hosting Provider
Now that you have your
Table 3: Qualitative results of caching performance. We generate 100 tokens for each prompt.